OpenStudy (anonymous):

What is the factored form of 9x2 + 12x + 4?

OpenStudy (anonymous):

(3x + 2)(3x+2)

OpenStudy (anonymous):

ok thank you(:

OpenStudy (anonymous):

you're welcome :)

hero (hero):

\[\space\space\space\space\space\space9x^2 + 12x + 4 \\= 9x^2 + 6x + 6x + 4 \\= 3x(3x + 2) + 2(3x + 2) \\= (3x + 2)(3x + 2) \\= (3x + 2)^2 \]